# All Entities are Not Created Equal: Examining the Long Tail for Ultra-Fine Entity Typing

### v2

@@ Appendix B Temporal dynamics of search API hits
We use the Google Search API to approximate the distribution of entity frequencies that models have seen during training. While convenient, this approach may potentially ignore the temporal changes that might occur in the distributions of such entities. This is especially important as models we discuss in our work (BERT, BART and LLAMA) have been trained at different points within the last decade. To ensure that these dynamics do not significantly impact our claims, we compare the correlation between the LM predictions and the API data capped at 2018 and 2024. We find results to be largely consistent across these two time periods (See Tab. 3). On further examination, we find that between 2018 and 2024 only 39 entities from the test set (<2%) change their bin classification. For our main results, we rank entities using the 2024 results.
@@
Bert-base-uncased 0.873 0.873
Bart-large 0.645 0.649
Llama-3.2-1B-Instruct 0.849 0.847
@@ Appendix C UFET test bin distribution
Figure 3: Entity distribution across UFET test bins
@@
Appendix D Prompt used to calculate the probability to recover an entity for LLAMA
The prompt used to calculate the probability to recover an entity for LLAMA is:
@@ Appendix D Prompt used to calculate the probability to recover an entity for LLAMA
Appendix E Method for Generating Masks and Calculating Entity Probability
LLAMA (Causal LM): Since LLAMA Dubey et al. (2024) is not pre-trained with a MLM objective, we reframe the task as a Fill-in-the-Blank problem using Llama-3.2-1B-Instruct. We provide the model with a prompt (See App. D) and compute the probabilities through a conditional generation process, one token at a time.
Appendix F Scatter plots of Entity Recovery Probability in BERT, BART and LLAMA against Search Engine API hits
(a) BERT: average probability vs. log search-engine hits
(b) LLAMA: average probability vs. search-engine hits
(c) BART: average probability vs. search-engine hits
Figure 5: Average UFET Entity Recovery Probabilty vs Search API Hits
@@ Appendix F Scatter plots of Entity Recovery Probability in BERT, BART and LLAMA against Search Engine API hits
Appendix G Effect of word splitting on entity recovery probability
To understand if the splitting of words negatively impacts the probability of recovering an entity as discussed in Sec. E, we plot the mean probability of entity recovery (see Fig. 6) against the number of tokens per word in an entity. While there is a marked dip in the recovery probability as we start encountering words being split, no clear trend emerges, suggesting that tokenizers alone cannot explain the dip in performance for certain entities. This suggest that our Search API method is a better, more nuanced proxy to approximate entity frequency as it pertains to parametric knowledge.
(a) BERT-Base-Uncased
(b) BART-Large
(c) LLAMA-3.2-1B-Instruct
Appendix H Models for entity typing
H.1 BERT - Baseline
@@
H.2 Llama3/Qwen3 - Baseline
@@
H.3 UFET-LSTM Choi et al. (2018)
@@
H.5 Box4Types Onoe et al. (2021)
Box4Types deploys box emeddings to effectively capture the hierarchies of types. The model represents both types and entity mentions as boxes. Each mention and its context are fed into a BERT-based model to embed that mention in our box space; essentially, the model leverages typological clues present in the surface text to hypothesize a type representation for the mention. This helps the model capture latent hierarchies better than the vector based counterparts. We recreated the original approach described in paper and evaluated it against our bins in Table 9.
H.6 Label Reasoning Network Liu et al. (2021)
Label Reasoning Network sequentially reasons about fine-grained entity labels by discovering and exploiting knowledge about label dependencies entailed by the data. These implicitly and explicitly entailed dependencies provide critical information which help the model overcome limitations of baseline LM approaches. The BERT based model leverages deductive and inductive reasoning. We recreated the models (without retrieval) as described in the paper and reported the results against our bins in table 8.
Appendix I Fine grained evaluation of the models studied
We look at the performance of the discussed models across bins and label granularities (Coarse, Fine, UltraFine) as first proposed by Choi et al. (2018).
@@ Appendix I Fine grained evaluation of the models studied
Overall Coarse Fine UltraFine
@@
Table 5: LLAMA 3.3 70B
Overall Coarse Fine UltraFine
Subset P(±σ) R(±σ) F1(±σ) P(±σ) R(±σ) F1(±σ) P(±σ) R(±σ) F1(±σ) P(±σ) R(±σ) F1(±σ)
@@
Table 6: Qwen3 32B
Overall Coarse Fine UltraFine
Subset P(±σ) R(±σ) F1(±σ) P(±σ) R(±σ) F1(±σ) P(±σ) R(±σ) F1(±σ) P(±σ) R(±σ) F1(±σ)
@@
Overall Coarse Fine UltraFine
@@
Overall Coarse Fine UltraFine
@@
Subset Overall Coarse Fine UltraFine
@@
Overall Coarse Fine UltraFine

### v3

@@ Appendix B Temporal dynamics of search API hits
We use the Google Search API to approximate the distribution of entity frequencies that models have seen during training. While convenient, this approach may potentially ignore the temporal changes that might occur in the distributions of such entities. This is especially important as models we discuss in our work (BERT, BART and Qwen3) have been trained at different points within the last decade. To ensure that these dynamics do not significantly impact our claims, we compare the correlation between the LM predictions and the API data capped at 2018 and 2024. We find results to be largely consistent across these two time periods (See Tab. 4). On further examination, we find that between 2018 and 2024 only 39 entities from the test set (<2%) change their bin classification. For our main results, we rank entities using the 2024 results.
@@
Bert-base-uncased 0.883 0.885
Bart-large 0.711 0.716
Qwen3-0.6B 0.895 0.897
@@ Appendix C UFET test bin distribution
Appendix D Method for Generating Masks and Calculating Entity Probability
Qwen3 (Causal LM): Since Qwen3 Yang et al. (2025) is not pre-trained with an MLM objective, we reframe the task as a Fill-in-the-Blank problem using Qwen3-0.6B. We provide the model with a prompt (See App. E) and compute the probabilities through a conditional generation process, one token at a time.
Appendix E Prompt used to calculate the probability to recover an entity for Qwen3
The prompt used to calculate the probability to recover an entity for Qwen3 is:
@@ Appendix E Prompt used to calculate the probability to recover an entity for Qwen3
Appendix F Scatter plots of Entity Recovery Probability in BERT, BART and Qwen3 against Search Engine API hits
(a) BERT-Base-Uncased: Average probability vs. Log search‐engine hits
(b) BART-Large: Average probability vs. Log search‐engine hits
(c) Qwen3-0.6B: Average probability vs. Log search‐engine hits
Figure 6: Average UFET Entity Recovery Probability vs Search API Hits
@@ Appendix F Scatter plots of Entity Recovery Probability in BERT, BART and Qwen3 against Search Engine API hits
Appendix G Models for entity typing
G.1 BERT - Baseline
@@
G.2 Llama3/Qwen3 - Baseline
@@
G.3 UFET-LSTM Choi et al. (2018)
@@
G.5 Box4Types Onoe et al. (2021)
Box4Types deploys box embeddings to effectively capture the hierarchies of types. The model represents both types and entity mentions as boxes. Each mention and its context are fed into a BERT-based model to embed that mention in our box space; essentially, the model leverages typological clues present in the surface text to hypothesize a type representation for the mention. This helps the model capture latent hierarchies better than the vector-based counterparts. We recreated the original approach described in paper and evaluated it against our bins in Tab. 10.
G.6 Label Reasoning Network Liu et al. (2021)
Label Reasoning Network sequentially reasons about fine-grained entity labels by discovering and exploiting knowledge about label dependencies entailed by the data. These implicitly and explicitly entailed dependencies provide critical information which help the model overcome limitations of baseline LM approaches. The BERT based model leverages deductive and inductive reasoning. We recreated the models (without retrieval) as described in the paper and reported the results against our bins in Tab. 9.
Appendix H Fine grained evaluation of the models studied
We look at the performance of the discussed models across bins and label granularities (Coarse, Fine, Ultra-fine) as first proposed by Choi et al. (2018).
@@ Appendix H Fine grained evaluation of the models studied
Overall Coarse Fine Ultra-fine
@@
Table 6: Llama3.3-70B
Overall Coarse Fine Ultra-fine
Subset P(±ς) R(±ς) F1(±ς) P(±ς) R(±ς) F1(±ς) P(±ς) R(±ς) F1(±ς) P(±ς) R(±ς) F1(±ς)
@@
Table 7: Qwen3-32B
Overall Coarse Fine Ultra-fine
Subset P(±ς) R(±ς) F1(±ς) P(±ς) R(±ς) F1(±ς) P(±ς) R(±ς) F1(±ς) P(±ς) R(±ς) F1(±ς)
@@
Overall Coarse Fine Ultra-fine
@@
Overall Coarse Fine Ultra-fine
@@
Subset Overall Coarse Fine Ultra-fine
@@
Overall Coarse Fine Ultra-fine